OpenStudy (anonymous):

Solve the absolute value inequality: |4x + 3| >= 19

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

can u show me the right stepss :)

OpenStudy (apoorvk):

Well, one simple funda: \(|t|\ge y\) \(=>t\le -y\) and \(t\ge y\) Can you use this funda to begin with?

OpenStudy (anonymous):

|4x+3|>=19 4x+3>=19 or 4x+3<=-19 x>=4 or x<=-22/4

OpenStudy (anonymous):

soo x<=5.5 or 4<x

OpenStudy (anonymous):

thanks chang chung

OpenStudy (anonymous):

u welcome